X Company has the following estimated costs for 2020:

Direct materials $55,200
Direct labor 21,000
Factory supplies 19,400
Factory supervision 31,500
Marketing research 45,600
Factory rent 48,200
Sales salaries 44,000
Factory insurance 64,800

X Company estimates that direct labor hours for the year will be 1,500, and machines will be operated for 69,000 hours. If overhead is allocated on the basis of direct labor cost, what will the 2020 overhead rate be?
Answer #1
Answer: $109.27 per direct labor hour

Overhead = 19,400+31,500+48,200+64,800 = 163,900

Direct labor hours = 1,500
since overhead is allocated based on direct labor hours, the rate is:
163,900/1,500 = 109.27
Answer: $109.27 per direct labor hour

Please like